OpenStudy (dls):

Find the number of solutions to the equation : x1+x2+x3+x4+x5 = 21 where xi is a non negative integer Provided that 1) 0<=x1<=10 2) x1 is from 0 to 3 , x2 is from 1 to 4, x3 >=15

OpenStudy (dls):

I know that the standard formula for calculating the number of non negative integral solutions is given by \(\Large C(n+r-1,r-1)\). And I know if there are constraints like \(\Large x_i > \alpha\) then we replace \(\Large x_i\) with \(\Large x_i - \alpha\) .(Hopefully I have these facts right). But I am not sure how to deal with these 2 conditions.

OpenStudy (dls):

@ganeshie8

ganeshie8 (ganeshie8):

Hey you may use generating functions

OpenStudy (dls):

Can you give a good place to learn that ? :/ my textbook has described it in a complicated way :|

ganeshie8 (ganeshie8):

I have a very good tutorial on generating functions, one sec

ganeshie8 (ganeshie8):

Here it is http://db.math.ust.hk/notes_download/elementary/algebra/ae_A11.pdf @DLS

ganeshie8 (ganeshie8):

1) 0<=x1<=10 you may proceed like this : since \(x_1\) can take values from \(0\) to \(10\), it can be 0 or 1 or 2 or .... or 10 the corresponding factor for \(x_1\) in the generating function is \((x^0+x^1+x^2+\cdots + x^{10})\)

ganeshie8 (ganeshie8):

There are no restrictions on other four variables, so the corresponding factor for each of them is \((x^0+x^1+x^2+\cdots)\)

ganeshie8 (ganeshie8):

The overall generating function is then given by \[G(x)= (x^0+x^1+x^2+\cdots + x^{10}) (x^0+x^1+x^2+\cdots )^4\]

ganeshie8 (ganeshie8):

The answer for part 1 is the coefficient of \(x^{21}\) of \(G(x)\)

ganeshie8 (ganeshie8):

Can you think of a way to find the coefficient of \(x^{21}\) of \(G(x)\) ?

OpenStudy (dls):

oh this thing is called generating functions? :D didn't know that. we can use multinomial theorem maybe?

ganeshie8 (ganeshie8):

we don't really need to get all that sophisticated

ganeshie8 (ganeshie8):

binomial theorem will do :)

OpenStudy (dls):

how are you planning to apply binomial here :o

ganeshie8 (ganeshie8):

\[G(x)= (x^0+x^1+x^2+\cdots + x^{10}) (x^0+x^1+x^2+\cdots )^4\] How are you planning to avoid binomial theorem here ?

ganeshie8 (ganeshie8):

\[G(x)= (x^0+x^1+x^2+\cdots + x^{10}) (x^0+x^1+x^2+\cdots )^4\\~\\ =(\dfrac{1-x^{11}}{1-x}) (\dfrac{1}{1-x})^4 \]

ganeshie8 (ganeshie8):

does that look okay to you?

ganeshie8 (ganeshie8):

I have just used the partial sum and infinite sum formulas of geometric series

OpenStudy (dls):

what did you do in the first term :o

ganeshie8 (ganeshie8):

\[x^0+x^1+x^2+\cdots + x^{10}\] Notice that this is a geometric series with first term = \(1\) and common ratio = \(x\)

ganeshie8 (ganeshie8):

remember the partial sum formula of geometric series ?

OpenStudy (dls):

oh right! sorry,my bad, yeah got that step!

ganeshie8 (ganeshie8):

good, try simplifying that expression a bit

OpenStudy (dls):

i can proceed now :)

ganeshie8 (ganeshie8):

familiar with extended binomial theorem for negative exponents ?

ganeshie8 (ganeshie8):

\[G(x)= (x^0+x^1+x^2+\cdots + x^{10}) (x^0+x^1+x^2+\cdots )^4\\~\\ =(\dfrac{1-x^{11}}{1-x}) (\dfrac{1}{1-x})^4\\~\\ = (1-x^{11})(1-x)^{-5} \]

ganeshie8 (ganeshie8):

see if that looks okay

OpenStudy (dls):

yep :)

OpenStudy (dls):

we can open the bracket now i guess

OpenStudy (dls):

\[=>(1-x^{11}) - x^{11}(1-x)^5\]

OpenStudy (dls):

we could apply binomial theorem on the 2nd term :o

OpenStudy (dls):

yeah that's fine too :)

ganeshie8 (ganeshie8):

\[=>(1-x^{11}) - x^{11}(1-x)^5\] how did you get that ?

OpenStudy (dls):

sorry the power is -5

ganeshie8 (ganeshie8):

\[G(x)= (x^0+x^1+x^2+\cdots + x^{10}) (x^0+x^1+x^2+\cdots )^4\\~\\ =(\dfrac{1-x^{11}}{1-x}) (\dfrac{1}{1-x})^4\\~\\ = (1-x^{11})(1-x)^{-5}\\~\\ =(1-x^{11} )(\sum\limits_{k=0}^{\infty}\dbinom{-5}{k} (-x)^k) \]

OpenStudy (dls):

I meant this..but it isn't required. \[\Large (1-x^{-5}) - x^{11}(1-x^{-5})\]

OpenStudy (dls):

I thought of writing it like that because the coefficient of x^17 would only come from the 2nd term.

ganeshie8 (ganeshie8):

do you mean I meant this..but it isn't required. \[\Large (1-x)^{-5} - x^{11}(1-x)^{-5}\] ?

OpenStudy (dls):

yeah :P equation editor is trolling me today :/

ganeshie8 (ganeshie8):

that looks god as you can see the first term \((1-x)^{-5}\) contains an \(x^{21}\) term and the second term \(x^{11}(1-x)^{-5}\) also contains an \(x^{21}\) term we need to find them both

ganeshie8 (ganeshie8):

whats the coefficient of \(x^{21}\) in the expansion of \((1-x)^{-5}\) ?

OpenStudy (dls):

\[\large C(-5,r)1^r (-x)^{-n-r} => C(-5+r-1,r)(-x)^{-5-r}\]

OpenStudy (dls):

-5-r = 21 r = -26?

OpenStudy (dls):

This is the 2nd or 3rd time I happen to use binomial theorem for negative index :|

ganeshie8 (ganeshie8):

you're making it way more complicated than it actually is

OpenStudy (dls):

lol sorry

ganeshie8 (ganeshie8):

whats the coefficient of \(x^{21}\) in the expansion of \((1-x)^{-5}\) ? the term is : \(C(-5, 21) (-x)^{21}\)

ganeshie8 (ganeshie8):

therefore the coefficient of \(x^{21}\) is simply \(-C(-5, 21)\)

ganeshie8 (ganeshie8):

How we evaluate that is another story

OpenStudy (dls):

ohh..all this is because I haven't got the grip of this topic anymore :/ I had done it in XIIth or so, due to the gap in first year I have totally forgotten it :/ sorry..but I happen to remember it now :)

ganeshie8 (ganeshie8):

its all good :) see if you can find the "required" coefficient in second term

ganeshie8 (ganeshie8):

\[\Large (1-x)^{-5} - x^{11}\color{red}{(1-x)^{-5}}\]

ganeshie8 (ganeshie8):

do we need the coefficient of \(x^{10}\) in that red term ?

OpenStudy (dls):

yep..we need all such indices which can add up with 11 to produce 21

ganeshie8 (ganeshie8):

there is only one number that produces 21 when added to 11

ganeshie8 (ganeshie8):

and that number is 10

ganeshie8 (ganeshie8):

once you see how all this works, it is a very simple problem actually...

OpenStudy (dls):

yeah hopefully this would help in clearing alot of questions :| anyway, how do we get to know which all powers will (1-x)^-5 have ?

ganeshie8 (ganeshie8):

\((1-x)^{-5} = \sum\limits_{k-0}^{\infty}C(-5, k)(-x)^k\)

ganeshie8 (ganeshie8):

as you can see it will have all the powers of x

ganeshie8 (ganeshie8):

we're just looking for the coefficient of \(x^{21}\)

OpenStudy (dls):

I don't know why I'm getting so much confused lol, I get it :D

ganeshie8 (ganeshie8):

its okay, read the first 3 pages of the pdf i gave you earlier

OpenStudy (dls):

ALRIGHT! so we have the coefficient from the 2nd term now.

OpenStudy (dls):

and the first term as well

OpenStudy (dls):

just a last step to evaluate that term

ganeshie8 (ganeshie8):

what is the coefficient from second term ?

OpenStudy (dls):

C(-5,10)

OpenStudy (dls):

+C(-5,21) from the first term

OpenStudy (dls):

= total answer

ganeshie8 (ganeshie8):

Excellent! careful about the signs though

OpenStudy (dls):

oh yes. there is a (-1)^n hanging around there

OpenStudy (dls):

only the first term should have it.

ganeshie8 (ganeshie8):

\[\Large (1-x)^{-5} - x^{11}\color{red}{(1-x)^{-5}}\] \(x^{21}\) coefficient from first term is \(-C(-5,21)\) \(x^{21}\) coefficient from second term is \(C(-5,10)\) subtract them to get \[-C(-5,21)-C(-5,10)\]

ganeshie8 (ganeshie8):

that is our final answer

OpenStudy (dls):

yep!!

ganeshie8 (ganeshie8):

evaluate that

ganeshie8 (ganeshie8):

treat that same as a regular binomial coefficient

ganeshie8 (ganeshie8):

\[C(n,k) = \dfrac{n(n-1)(n-2)\cdots (n-k+1)}{k!}\]

ganeshie8 (ganeshie8):

\[C(-5,21) = \dfrac{-5(-5-1)(-5-2)\cdots (-5-21+1)}{21!}\]

ganeshie8 (ganeshie8):

whatever you get

ganeshie8 (ganeshie8):

it should simplify nicely though

ganeshie8 (ganeshie8):

lets try

ganeshie8 (ganeshie8):

\[C(-5,21) = \dfrac{-5(-5-1)(-5-2)\cdots (-5-21+1)}{21!}\\~\\ =\dfrac{(-1)^{21}5(5+1)(5+2)\cdots (5+21-1)}{21!} \]

ganeshie8 (ganeshie8):

Oh my! that is same as \[C(-5,21) = \dfrac{-5(-5-1)(-5-2)\cdots (-5-21+1)}{21!}\\~\\ =\dfrac{(-1)^{21}5(5+1)(5+2)\cdots (5+21-1)}{21!}\\~\\ =(-1)^{21}C(5+21-1,~21) \]

ganeshie8 (ganeshie8):

that is just a regular binomial coefficient which you can evaluate using ur calculator

ganeshie8 (ganeshie8):

That is so awesome! In general do we have below formula to convert negative binomial coefficients to positive : \[C(-n,~r) = (-1)^r C(n+r-1~r)\]

ganeshie8 (ganeshie8):

?

OpenStudy (dls):

ohh..nice :D so did you generalize a formula for negative coefficient btw ? \[\Large (-1)^r C(r-n+1)\]

ganeshie8 (ganeshie8):

Now I see... you have tried to use that formula earlier right ?

OpenStudy (dls):

haha yes :p

ganeshie8 (ganeshie8):

Oh my! that is same as \[C(-5,21) = \dfrac{-5(-5-1)(-5-2)\cdots (-5-21+1)}{21!}\\~\\ =\dfrac{(-1)^{21}5(5+1)(5+2)\cdots (5+21-1)}{21!}\\~\\ =(-1)^{21}C(5+21-1,~21)\\~\\ =-C(25,~21) \]

ganeshie8 (ganeshie8):

does that look good ?

OpenStudy (dls):

yeah alot better :D

ganeshie8 (ganeshie8):

wolfram says the final answer is 11649 http://www.wolframalpha.com/input/?i=-%28-5+choose+21%29-%28-5+choose+10%29

ganeshie8 (ganeshie8):

il let you finish it off :)